OpenStudy (anonymous):

2(x-3)=3(x+7), solve, must show work.

OpenStudy (anonymous):

also which side do i start on?

OpenStudy (anonymous):

2(x-3)=3(x+7) First distribute 2 to x - 3 and 3 to x + 7 2x - 6 = 2x + 21

OpenStudy (anonymous):

2(x-3)=3(x+7) 2x - 2*3 = 3x + 3*7 2x - 6 = 3x + 21 2x -6 -3x = 21 2x - 3x = 21 + 6 -x = 27 x = -27

OpenStudy (ammarah):